Determine the surface area of the curb. Do not include the area of the ends in the calculation. 30 4 m 150 mm 150 mm 100 mm 150 mm

1
Expert's answer
2021-11-25T01:41:42-0500

Calculate the arc length at the mean position of the curb.

"L=r \theta\\
=(4+0.15)(30* \frac{\pi}{180}\\
=2.173 m"

Calculate the volume of the curb.

"V= A_c *L\\
=(0.3*0.15+ \frac{1}{2}*0.15*0.1)*2.173\\
=0.114 m^3"

Calculate the surface area excluding the area of the ends.

"A_s= Perimeter \times L\\
= (0.25+0.3+0.15+0.15+ \sqrt{0.15^2+0.1^2})*2.173\\
=2.238 m^2"